Accelerator Update for October 20, 2008 - October 22, 2008
Monday Oct 20
The day shift began with Operations monitoring the Tevatron (TeV) store 6501 (50.76E30) and the Recycler stash (341.7E10), with the Antiproton Source (Pbar) stacking (20.9E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE, NuMI, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
At 8:21 AM, a Booster RF station (BRF15) tripped off. It reset.
Operations put Pbar into standby mode at 8:32 AM.
At 8:55 AM, Operations reported that the Klystron Vernier had tripped off eight times this morning on a LCW flow fault. It reset.
A RF expert reported at 9:08 AM that he was changing the modulator for a MI RF station (MIRF4).
Linac experts reported at 9:55 AM that they had replaced the flow turbine for the Klystron Vernier.
At 10:03 AM, the Booster Chopper (B:CHOP) tripped off. Experts investigated.
At 10:13 AM, the MI Power Supply interlock LOOP (MIPSLOOP) tripped off due to a volts-to-ground indication on the MI-40U power supply. Operations talked with an expert and then reset everything.
At 10:45 AM, the Pelletron tripped off. An expert investigated and reported that the chain power supply current was too high. He had the Pelletron back up and running about 13 minutes later.
Linac experts reported at 12:26 PM that they had brought KRF7 online, but the Linac HBENDS power supply had tripped off. Operators reset the supply.
Operations established beam to MiniBooNE at 12:28 PM.
Operations established beam to Pbar and NuMI at 12:36 PM.
Power supply experts reported at 12:36 PM that they had repaired the MI50U power supply and brought it back online.
Operations terminated store 6501 at 1:17 PM.
Operations put the TeV into shot setup at 1:28 PM.
At 1:40 PM, a MI Bend magnet power supply began having regulation problems. Over a five-minute period this caused about a 4.5E10 loss of antiprotons from the stash.
At 1:42 PM, the MIPSLOOP tripped off for the first of two times in one minute on a filter choke indication on the MI-40 Upper and Lower power supplies.
Water group experts reported at 3:20 PM that they had made up water for the MI LCW system.
Operations established store 6502 at 3:38 PM with an initial luminosity of 276.97E30.
A MI expert reported at 3:46 PM that he was beginning a multiwire study in the MI-8 beam transfer line.
Operations disabled the MTest beam at 6:01 PM.
At 8:39 PM, Operations reported that a Linac quadrupole power supply’s (QPS203) output had dropped to zero. Operators swapped out its control card with QPS205.
At 9:25 PM, KRF5 tripped off. It reset.
At 11:51 PM, BRF7 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed antiprotons six times during the evening shift.
Tuesday Oct 21
The midnight shift began with Operations monitoring store 6502 (105.43E30) and the stash (202.7E10), with Pbar stacking (27.2E10), and with the MiniBooNE and NuMI experiments taking beam.
At 3:26 AM, Operations reported that the temperature for the I- and H- Source domes rose by about five degrees. Operations contacted the Duty Mechanic who later reported that he had reset the air conditioning unit.
At 5:05 AM, Operations reported that after the Duty Mechanic had reset the AC unit, the temperature for the H- Source dropped, but the I- Source continued to climb.
At 5:49 AM, MIRF5 tripped off and would not reset remotely. It did reset locally.
Operations established beam to MTest at 6:39 AM.
At 6:56 AM, BRF16 tripped off. It reset.
Operations reported at 7:52 AM that the Recycler had stashed seven times during the midnight shift.
At 8:18 AM, MIRF7 tripped off. It reset.
Interlock experts began a Meson MTest safety system test at 8:32 AM.
Operations terminated store 6502 at 10:02 AM.
At 10:14 AM, LRF5 tripped off. It reset.
Operations put the TeV into shot setup at 10:24 AM.
Booster experts investigated and reported at 11:27 AM that the 400 MeV area trips were due to arcing in the Booster Chopper off tube. They made repairs and Pbar resumed stacking.
At 1:18 PM, BRF7 tripped off for the first of two times in seven minutes. It reset.
Operations prepared the Pelletron for a one-hour access at 1:29 PM.
Operations established store 6503 at 1:45 PM with an initial luminosity of 289.69E30.
At 2:25 PM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 & 8 to also trip off. Everything reset.
At 2:48 PM, Operations reported that NuMI had tripped off twice due to MI RF problems.
TeV experts began a 20-minute optics study at 4:02 PM.
Operations reported at 4:06 PM that the Recycler had stashed four times during the day shift.
Operations disabled the MTest beam at 6:22 PM.
A power supply expert reported at 7:09 PM that he had repaired and dried out D:IB.
Pbar resumed stacking at 7:40 PM.
Operations reported at 9:21 PM that earlier the Cryo Coordinator was called in to investigate the Pbar cryo system. The system had warmed slightly over the last couple of days. He reported that cryo technicians tightened some belts on the Pbar dry engine.
At 11:17 PM, LRF3 tripped off. It reset.
At 11:59 PM, KRF7 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Wednesday Oct 22
The midnight shift began with Operations monitoring store 6503 (86.18E30) and the stash (176.8E10), with Pbar stacking (35.6E10), and with the MiniBooNE and NuMI experiments taking beam.
At 1:11 AM, KRF5 tripped off. It reset.
At 1:17 AM, MIRF7 tripped off. It reset.
At 2:49 AM, Operations reported that the belt on the Pbar dry engine was slipping. Operations contacted the Cryo Coordinator and experts. They reported that the flywheel needed to be replaced.
A TeV expert began a Tevatron Electron Lens (TEL2) beam-beam compensation study at 5:56 AM.
Operations reported at 6:09 AM that cryo system experts began replacing the Pbar dry engine flywheel.
Operations established beam to MTest at 6:20 AM.
Operations reported at 7:55 AM that the Recycler had stashed seven times during the midnight shift.
Cryo system experts reported at 7:56 AM that they had completed the Pbar dry engine work.
The Plans for Wednesday
The plans for today are to allow a Linac access, after the next store goes into the TeV, for MTA - MuCool work. The Run Coordinator will allow other machine work to occur during Linac access.
Comments and Suggestions
|last modified 10/22/2008 email Fermilab|
|Security, Privacy, Legal|